# Automatically Advancing LLM Expertise in Technology Judgment

### v3

@@ Appendix B Related Literature
Relations to other benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding. But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after [McIntosh et al., 2024, Wu et al., 2024, Davis, 2023]. High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension. In particular, current evaluation datasets frequently allow models to exploit statistical associations or training data overlaps to get the right answers without genuine understanding  [^5] https://www.sciencenews.org/article/ai-understanding-reasoning-skill-assess. Even when test content is novel, models can exploit formatting quirks. Alzahrani et al. [Alzahrani et al., 2024] showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions. Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application. Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics [Arakelyan et al., 2024]. Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities [Burnell et al., 2023]. In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities. Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase [Asthana et al., 2024] or contrast and counterfactual sets [Lewis and Mitchell, 2024]  [^6] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content [West et al., 2024]..
@@
Organizing internal and external knowledge: A core challenge in LLM understanding is how to activate the vast latent knowledge embedded in the model’s billions of parameters. Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand [Borgeaud et al., 2022, Hoang et al., , Hoang et al., 2023, Jaiswal et al., 2024, Pan et al., 2025]. Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted. The model might skillfully "reason" with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge. This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial [Schulhoff et al., 2024, Gopnik and Wellman, 2012]. To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources. Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence [Prince et al., 2024]. Lewis et al. [Lewis et al., 2020] demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks. Borgeaud et al. [Borgeaud et al., 2022] introduced the RETRO (Retrieval-Enhanced Transformer) model. A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval. He et al. [He et al., 2022] showed that an LLM with a retrieval step produces more faithful, grounded understanding – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness. In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries. This is likely because LLMs gain new information during retrieval. LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data [Shumailov et al., 2024]. By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarities and differences.
@@
Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity [He et al., 2020, Gao et al., 2021, Chen et al., 2020, Cohan et al., 2020]. Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure. While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.
Self-questioning and self-growth of LLMs: LLMs have recently been prompted to engage in self-questioning and introspection to improve their reasoning and conceptual understanding [Huang et al., 2023, Qu et al., 2024, Song et al., 2024, Wang et al., 2023c, b]. Instead of only generating direct answers, many contemporary models are encouraged to ask themselves clarifying questions and work through sub-problems, mimicking a Socratic or pedagogical approach. Early prompting techniques like chain-of-thought (CoT) [Wei et al., 2022] demonstrated that guiding LLMs to reason step-by-step markedly boosts performance on complex tasks. Building on this, researchers introduced methods for the model to generate its own intermediate questions. For example, self-ask prompting [Press et al., 2022] explicitly guides the model to pose follow-up questions to itself and then answer them in sequence, effectively decomposing complex queries into tractable pieces. Beyond decomposition, recursive self-reflection techniques have emerged to let LLMs iteratively improve their answers. Madaan et al. [Madaan et al., 2023] introduce Self-Refine, a framework where an LLM generates an initial answer and then critiques and revises its own output in multiple rounds. Shinn et al. [Shinn et al., 2023] propose Reflexion, in which an agent-model “verbally reflects” on feedback from previous trials and stores these self-generated insights in memory to inform subsequent attempts. Huang et al. [Huang et al., 2023] showed that even without external labels, an LLM can “self-train” by generating reasoning traces for unsolved problems and fine-tuning on them, yielding notable gains in mathematical and commonsense tasks. Miao et al. present SelfCheck [Miao et al., 2024], a zero-shot verification schema enabling LLMs to assess their own step-by-step reasoning without external resources, thereby enhancing accuracy in complex tasks through internal consistency checks. Similarly, techniques like least-to-most prompting [Zhou et al., 2022] encourage the model to teach itself gradually, which has improved performance on multi-step reasoning challenges. Other work has explored letting models self-evaluate or re-rank their answers (e.g., self-consistency voting [Wang et al., 2023a]) and prompts them to explicitly verify or justify their solutions, analogous to a student checking their work. These introspective or pedagogically-inspired strategies are motivated by human learning: students often learn better by asking themselves questions and reflecting on mistakes. Empirically, such approaches can bolster an LLM’s conceptual grasp and reduce reasoning errors. Nevertheless, there are also limitations and mixed findings. It is hard to evaluate the quality of self-generated questions because LLMs may not be able to answer their self-generated questions [Balepur et al., 2025]. As we demonstrate, large models are much more likely to productively answer small model questions than the converse. Not all tasks benefit uniformly from self-reflection. Li et al. [Li et al., 2024] find that prompting GPT-4 to reflect on its answers helps on a truth-focused QA task but hinders performance on a multi-hop knowledge task. Thus, while self-questioning and introspective prompts represent a powerful paradigm for enhancing LLM reasoning, careful tuning is needed to ensure the model’s “inner dialogue” remains productive and does not overthink or hallucinate.
@@
(b) Few-shot CoT helps identify similarities across patent abstracts and their rewritten versions.
Figure 6: Few-shot CoT reasoning can distinguish the rephrased abstracts (rewritten by the GPT-o3 model), but cannot distinguish their counterpart (another unique patent).
@@
Following on prior sections that focus on distinguishing deceptively similar patents, in this section, we explore whether and how LLMs are able to merge identical but paraphrased patent ideas. Our approach to measuring LLMs’ ability to generalize and identify the same ideas across distinct articulations is to generate rephrased versions of each patent abstract using the OpenAI ChatGPT o3 model with the prompt “rephrase the patent, but keep its meaning.” We then task the LLaMA-3.1-8B-instruct model with identifying whether a patent abstract and its paraphrase describe the same invention. This was designed to parallel the original distinguishing experiment, in which we randomly sampled 1,000 patents and identified pairs that were most semantically similar. We then asked the LLaMA-3.1-8B-instruct model to determine whether the two abstracts described the same invention, using a human-annotated ground truth indicating that they were different. In raw (zero-shot) judgments, the model performed poorly on both merging and splitting tasks. We then introduced zero-shot CoT prompting (“think step-by-step”) and few-shot CoT, where we supplied examples of both merging (paraphrased) and splitting (semantically similar but distinct) patent pairs. Few-shot CoT substantially improved the model’s accuracy in recognizing paraphrased abstracts as equivalent—but had minimal effect on its ability to distinguish between closely related yet different inventions.
@@
This pattern suggests that CoT reasoning helps the model overcome basic semantic similarity checks and identify paraphrases, likely by encouraging step-by-step alignment at the surface level. Our hypothesis is that the model passes the rephrasing test primarily through pattern matching and semantic overlap, rather than through true conceptual integration. This explains why CoT improves performance on the rephrasing task: the test is relatively easy in the sense that our rewritten version keeps semantic similarity high, such that shallow strategies suffice. In contrast, differentiating between semantically similar but substantively different patents remains hard, even with CoT, because it requires more than aligned phrasing—it demands conceptual depth.
@@
In principle, a more advanced rephrasing model—one that intentionally avoids superficial alignment and instead rewrites concepts in deeper, structure-preserving ways—might push the LLM closer to the more difficult task of distinguishing near matches. Designing such a rephrasing system is far from trivial. Thus, while our current rewrite approach does not fully obscure surface similarity, the model’s success only when the task is easy reinforces the motivation behind our benchmark: LLMs rely heavily on surface-level semantic similarity, rather than demonstrating a deeper understanding of the underlying concepts. In this way, our dataset represents a frontier capability for all contemporary LLMs, even though larger models outperform smaller ones.
@@ Appendix D The Patent Dataset
The patent dataset comprises 1,319,184 patent pairs in total, drawn from five Cooperative Patent Classification (CPC) classes: H05 – Electric techniques not otherwise provided for; H04 – Electric communication technique; G06 – Computing; Calculating; Counting; H03 – Basic electronic circuitry; and G11 – Information storage. Within these, the patents are further categorized into 41 CPC subclasses: [’H05B’, ’H04L’, ’H05H’, ’H05K’, ’G06V’, ’G06N’, ’G06T’, ’H04N’, ’G06F’, ’G06Q’, ’H03M’, ’G06K’, ’H04B’, ’H04R’, ’G11B’, ’G11C’, ’H04M’, ’H03B’, ’H03F’, ’H03G’, ’H03H’, ’H03K’, ’H03L’, ’H04H’, ’H04J’, ’H04K’, ’H04W’, ’H04Q’, ’H04S’, ’H05F’, ’H05G’, ’H03C’, ’H03J’, ’G06E’, ’H03D’, ’G06G’, ’G06M’, ’G06C’, ’H05C’, ’G06J’, ’G06D’]. Out of all matched patent pairs in our dataset, 82.55% share the same CPC class, and 68.10% share the same, more specific CPC subclass. In contrast, only 10.50% of the pairs were granted in the same year.
This is an example patent pair where all LLaMA 3.1 instruct models — 8B, 70B, and 405B - fail to distinguish. We also show a scientific paper abstract to help readers perceive the distinct writing styles between patents and papers.
Patent 1: A wireless device (18) is configured for use in a wireless communication system. The wireless device (18) is configured to receive, from a radio network node (12), a multicast or broadcast transmission (14) and downlink control information indicating a number of repetitions with which the multicast or broadcast transmission (14) is transmitted. The wireless device (18) is also configured to operate according to a negative acknowledgement only feedback scheme in which the wireless device (18) is configured to transmit a negative acknowledgement to the radio network node (12) if decoding of the multicast or broadcast transmission (14) fails using the indicated number of repetitions and to refrain from transmitting a positive acknowledgement to the radio network node (12) if decoding of the multicast or broadcast transmission (14) succeeds.
@@
An Arxiv paper including the answer from the sampled patent pair: The Third Generation Partnership Project (3GPP) has recently published its Release 16 that includes the first Vehicleto-Everything (V2X) standard based on the 5G New Radio (NR) air interface. 5G NR V2X introduces advanced functionalities on top of the 5G NR air interface to support connected and automated driving use cases with stringent requirements. This paper presents an in-depth tutorial of the 3GPP Release 16 5G NR V2X standard for V2X communications, with a particular focus on the sidelink, since it is the most significant part of 5G NR V2X. The main part of the paper is an in-depth treatment of the key aspects of 5G NR V2X: the physical layer, the resource allocation, the quality of service management, the enhancements introduced to the Uu interface and the mobility management for V2N (Vehicle to Network) communications, as well as the co-existence mechanisms between 5G NR V2X and LTE V2X. We also review the use cases, the system architecture, and describe the evaluation methodology and simulation assumptions for 5G NR V2X. Finally, we provide an outlook on possible 5G NR V2X enhancements, including those identified within Release 17.
@@ Appendix E Prompts
Prompt for Question Generation System Prompt: You are an expert in patent comprehension. Your task is to generate structured questions that assess a reader’s background knowledge necessary to understand a given patent abstract. The questions should focus on foundational concepts, principles, and applications relevant to the patent’s domain without explicitly referencing the patent itself. Ensure the questions are framed generally and test for domain knowledge rather than details specific to the patent. The questions should follow Bloom’s Taxonomy and be categorized into three levels: 1. Remembering: Questions that assess the reader’s ability to recall key technical concepts, definitions, and fundamental principles relevant to the patent’s domain. These should focus on identifying and defining key terms, recognizing fundamental components, and understanding core principles of related technologies. 2. Understanding: Questions that assess the reader’s ability to explain how different elements of similar technologies function and interact. These should encourage summarization, interpretation, and explanation of related technical concepts and their roles in broader systems. 3. Applying: Questions that evaluate the reader’s ability to apply their knowledge by solving problems, making predictions, or considering real-world applications of the broader technological principles underlying the invention. These questions should be scenario-based and encourage practical thinking. Do not directly reference the patent abstract in any question. Instead, ensure all questions test knowledge that would be useful for understanding the patent without explicitly addressing its content. Maintain a structured and logical flow from basic recall to deeper conceptual understanding and practical application. Basic Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Abstract: {patent_summary} Generate a set of {qnum} questions that test a reader’s ability to recall fundamental concepts, key terms, and basic components necessary to understand the given patent abstract. The questions should focus on defining terms, identifying parts, and recognizing the primary function of the described technology. Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }} Conceptual Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Abstract: {patent_summary} Generate a set of {qnum} questions that assess a reader’s comprehension of the given patent abstract by requiring them to explain how different elements of the invention work together. The questions should encourage the reader to summarize the design, describe interactions between components, and interpret the intended improvements of the invention. Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }}
@@
Prompt for Answering Questions with Scientific Knowledge You are a scientific reasoning assistant. Use the following context to answer the question. Focus only on the information provided. Context: // retrieved scientific knowledge {context1} {context2} {context3} Question: {question} 1. Identify and summarize key points from the context relevant to the question. 2. Synthesize those points with your knowledge and try your best to answer the question. Return your answer in the following JSON format: {{ “answer”: “your answer” //your answer to the question }}
@@
Prompt for the Splitting Task Judgment You are a patent judge. Below are two patent abstracts with extra information. This extra information may or may not be useful in assessing the patents. You may choose to use it at your discretion. // This is for placebo answers Your task: judge whether they describe the same patent. Abstract 1: {abstract1} {qa_block1} Abstract 2: {abstract2} {qa_block2} Strictly return your concise answer in the following JSON format: {{ “label”: 1 // 1 if they describe the same patent, 0 if not “confidence”: 10 // your confidence score from 0 to 10, with 10 representing the highest confidence }} // Note: // This applies to the reverse questioning scenario (the major experiments in the paper): // "Do you think they are the same patent?" — when, in fact, the patents are judged to be *different* by human experts. // Invert this setup: Replace "same" with "different", and return 1 if the LLM concludes that the patents are *different*. This yields an alternative version of the experiment (See Section 3 Figure 2 in the paper).

### v4

@@ Appendix B Related Literature
Relations to current benchmarks: LLMs have achieved impressive scores on many NLP benchmarks, often giving the illusion of human-level reasoning or understanding. But researchers caution that excelling at a benchmark is not equivalent to possessing the general ability it is named after (McIntosh et al., 2024; Wu et al., 2024; Davis, 2023). High performance may reflect superficial pattern matching or memorization rather than true conceptual comprehension. In particular, current evaluation datasets frequently allow models to exploit statistical associations or training data overlaps to get the right answers without genuine understanding  [^6] https://www.sciencenews.org/article/ai-understanding-reasoning-skill-assess. Even when test content is novel, models can exploit formatting quirks. Current benchmark agreement tests are strongly influenced by benchmark-orthogonal factors such as question format (Wu et al., 2025). Alzahrani et al. (Alzahrani et al., 2024) showed that minor changes to the evaluation format – like shuffling multiple-choice answer order – can significantly alter MMLU performance and even flip model leaderboard rankings by several positions. Such brittleness indicates models are learning task-specific tricks (e.g., defaulting to a particular choice position or prompt wording) rather than demonstrating robust knowledge application. Minor wording tweaks could cause inconsistent outputs, suggesting that models are latching onto spurious lexical cues instead of grasping underlying semantics (Arakelyan et al., 2024). Aggregate metrics in existing benchmarks often obfuscate key information about where models tend to succeed or fail, broadcasting an inflated view of LLM capabilities (Burnell et al., 2023). In sum, many current benchmarks are static and shortcut-prone, failing to truly assess whether an AI understands concepts or is simply leveraging surface regularities. Recognizing these deficiencies, researchers have proposed more diagnostic evaluations to probe the conceptual depth of LLMs, such as contrastive evaluation using paraphrase (Asthana et al., 2024) or contrast and counterfactual sets (Lewis and Mitchell, 2024)  [^7] Some have expressed concern that even discriminative capacity falls prey to the “Generative AI Paradox”, wherein models can produce fluent, even expert-level outputs that far exceed their actual comprehension of the content (West et al., 2024)..
@@
Organizing internal and external knowledge: A core challenge in LLM understanding is how to activate the vast latent knowledge embedded in the model’s billions of parameters. Modern LLMs are trained on enormous corpora and ostensibly “know” a great deal, but this knowledge is stored in a highly compressed, distributed form not always accessible on demand (Borgeaud et al., 2022; Hoang et al., ; Hoang et al., 2023; Jaiswal et al., 2024; Pan et al., 2025). Indeed, there is a tension: an LLM may contain the facts needed for a task but still fail to recall them or apply them correctly when prompted. The model might skillfully "reason" with the information it does recall, yet still hallucinate a critical missing fact or overlook a subtle piece of world knowledge. This limitation is evident in tasks requiring niche or up-to-date information: the needed knowledge might exist somewhere in the model’s parameters, but eliciting it reliably through prompts is non-trivial (Schulhoff et al., 2024; Gopnik and Wellman, 2012). To address these issues, researchers have proposed retrieval-augmented generation (RAG) techniques, which marry LLMs with external knowledge sources. Rather than relying solely on the model’s compressed memory, a retrieval module fetches relevant documents (e.g., from Wikipedia or scientific literature) to provide the model with explicit evidence (Prince et al., 2024). Lewis et al. (Lewis et al., 2020) demonstrated the promise of this approach by generating more specific, diverse and factual language than a state-of-the-art parametric-only seq2seq baseline in NLP tasks. Borgeaud et al. (Borgeaud et al., 2022) introduced the RETRO (Retrieval-Enhanced Transformer) model. A RETRO model of only 7 billion parameters matched the performance of a 178B parameter model without retrieval. He et al. (He et al., 2022) showed that an LLM with a retrieval step produces more faithful, grounded understanding – their “rethinking with retrieval” method led to answers that stayed consistent with retrieved evidence, improving transparency and correctness. In general, augmenting LLMs with retrieval has been found to boost factual accuracy, reduce hallucinations, and expand the range of answerable queries. LLMs compress information, and so lose rare information in the “tail” of the distribution, leading to phenomena like model collapse when trained on recursively generated data (Shumailov et al., 2024). By searching through uncompressed information with RAG, LLMs also access critical and potentially rare information for making detailed decisions about complex concept similarities and differences.
@@
Contrastive learning: Our measures are also related to recent work on contrastive learning, which aims to improve model representations by explicitly training them to recognize semantic similarity and dissimilarity (He et al., 2020; Gao et al., 2021; Chen et al., 2020; Cohan et al., 2020). Contrastive methods work by pulling representations of semantically related inputs closer together while pushing those of unrelated inputs apart, fostering a latent space that reflects nuanced conceptual structure. While our evaluation does not involve contrastive training per se, it is motivated by a similar principle: if a model can consistently discern between distinct concepts, it demonstrates a deeper form of understanding—one likely to translate into stronger performance on downstream tasks.
@@
(b) Few-shot CoT helps identify similarities across patent texts and their rewritten versions.
Figure 5: Few-shot CoT reasoning can distinguish the rephrased patent texts (rewritten by the GPT-o3 model), as shown in the right panel (lower misjudge rate), but cannot distinguish their counterpart (another unique patent, as shown in the left panel). Y-axis always represents the response of "yes", i.e., the error rate (Left: yes when asking "are they the same" for close-but-different patent pairs. Right: yes when asking "are they different" for rephrased patents.)
@@
Following on prior sections that focus on distinguishing deceptively similar patents (the splitting task), in this section, we explore whether and how LLMs are able to merge identical but paraphrased patent ideas (the merging task). Our approach to measuring LLMs’ ability to generalize and identify the same ideas across distinct articulations is to generate rephrased versions of each patent text using the OpenAI ChatGPT o3 model with the prompt “rephrase the text, but keep its meaning.” We then task the LLaMA-3.1-8B-instruct model with identifying whether a patent text and its paraphrase describe different interventions (the correct answer: no). In raw (zero-shot) judgments, the model performed poorly on both merging and splitting tasks. We then introduced zero-shot Chain-of-Thought (CoT) prompting (“think step-by-step”) and few-shot CoT, where we supplied examples of both merging (paraphrased) and splitting (semantically similar but distinct) patent pairs. Few-shot CoT substantially improved the model’s accuracy in recognizing paraphrased texts as equivalent—but had minimal effect on its ability to distinguish between closely related yet different inventions.
@@
This pattern suggests that more reasoning helps the model overcome basic semantic similarity checks and identify paraphrases, likely by encouraging step-by-step alignment at the surface level. Our hypothesis is that the model passes the rephrasing test primarily through pattern matching and semantic overlap, rather than through true conceptual integration. The merging test is relatively easy in the sense that our rewritten version keeps key technical terms unchanged but only changes how the sentence is structured, such that shallow strategies suffice. In contrast, differentiating between semantically similar but substantively different patents remains hard, because it requires more than aligned phrasing—it demands conceptual depth.
@@
In principle, a more advanced rephrasing model—one that intentionally avoids superficial alignment and instead rewrites concepts in deeper, functionally equivalent ways—might push the LLM closer to the more difficult task of distinguishing near matches. Designing such a rephrasing system is far from trivial. Thus, while our current rewrite approach does not fully obscure surface similarity, the model’s success only when the task is easy reinforces the motivation behind our task: LLMs rely heavily on surface-level semantic similarity, rather than demonstrating a deeper understanding of the underlying concepts.
@@ Appendix D The Patent Dataset
The computer science patent dataset comprises 1,319,184 patent pairs in total, drawn from five Cooperative Patent Classification (CPC) classes: H05 – Electric techniques not otherwise provided for; H04 – Electric communication technique; G06 – Computing; Calculating; Counting; H03 – Basic electronic circuitry; and G11 – Information storage. Within these, the patents are further categorized into 41 CPC subclasses: [’H05B’, ’H04L’, ’H05H’, ’H05K’, ’G06V’, ’G06N’, ’G06T’, ’H04N’, ’G06F’, ’G06Q’, ’H03M’, ’G06K’, ’H04B’, ’H04R’, ’G11B’, ’G11C’, ’H04M’, ’H03B’, ’H03F’, ’H03G’, ’H03H’, ’H03K’, ’H03L’, ’H04H’, ’H04J’, ’H04K’, ’H04W’, ’H04Q’, ’H04S’, ’H05F’, ’H05G’, ’H03C’, ’H03J’, ’G06E’, ’H03D’, ’G06G’, ’G06M’, ’G06C’, ’H05C’, ’G06J’, ’G06D’]. The 177,903 biomedical patents are drawn from USPTO classes C01, C07–C14, and C21–C30. Out of all matched patent pairs in our dataset, 82.55% share the same CPC class, and 68.10% share the same, more specific CPC subclass. In contrast, only 10.50% of the pairs were granted in the same year.
This is an example patent pair where all LLaMA 3.1 instruct models — 8B, 70B, and 405B - fail to distinguish. We also show a scientific paper abstract to help readers perceive the distinct writing styles between patents and papers. Reference indexes are removed.
Patent 1: A wireless device is configured for use in a wireless communication system. The wireless device is configured to receive, from a radio network node, a multicast or broadcast transmission and downlink control information indicating a number of repetitions with which the multicast or broadcast transmission is transmitted. The wireless device is also configured to operate according to a negative acknowledgement only feedback scheme in which the wireless device is configured to transmit a negative acknowledgement to the radio network node if decoding of the multicast or broadcast transmission fails using the indicated number of repetitions and to refrain from transmitting a positive acknowledgement to the radio network node if decoding of the multicast or broadcast transmission succeeds.
@@
An Arxiv paper including the answer to questions about judging the sampled patent pair: The Third Generation Partnership Project (3GPP) has recently published its Release 16 that includes the first Vehicleto-Everything (V2X) standard based on the 5G New Radio (NR) air interface. 5G NR V2X introduces advanced functionalities on top of the 5G NR air interface to support connected and automated driving use cases with stringent requirements. This paper presents an in-depth tutorial of the 3GPP Release 16 5G NR V2X standard for V2X communications, with a particular focus on the sidelink, since it is the most significant part of 5G NR V2X. The main part of the paper is an in-depth treatment of the key aspects of 5G NR V2X: the physical layer, the resource allocation, the quality of service management, the enhancements introduced to the Uu interface and the mobility management for V2N (Vehicle to Network) communications, as well as the co-existence mechanisms between 5G NR V2X and LTE V2X. We also review the use cases, the system architecture, and describe the evaluation methodology and simulation assumptions for 5G NR V2X. Finally, we provide an outlook on possible 5G NR V2X enhancements, including those identified within Release 17.
@@ Appendix E Prompts
Prompt for Question Generation System Prompt: You are an expert in patent comprehension. Your task is to generate structured questions that assess a reader’s background knowledge necessary to understand a given patent. The questions should focus on foundational concepts, principles, and applications relevant to the patent’s domain without explicitly referencing the patent itself. Ensure the questions are framed generally and test for domain knowledge rather than details specific to the patent. The questions should follow Bloom’s Taxonomy and be categorized into three levels: 1. Remembering: Questions that assess the reader’s ability to recall key technical concepts, definitions, and fundamental principles relevant to the patent’s domain. These should focus on identifying and defining key terms, recognizing fundamental components, and understanding core principles of related technologies. 2. Understanding: Questions that assess the reader’s ability to explain how different elements of similar technologies function and interact. These should encourage summarization, interpretation, and explanation of related technical concepts and their roles in broader systems. 3. Applying: Questions that evaluate the reader’s ability to apply their knowledge by solving problems, making predictions, or considering real-world applications of the broader technological principles underlying the invention. These questions should be scenario-based and encourage practical thinking. Do not directly reference the patent text in any question. Instead, ensure all questions test knowledge that would be useful for understanding the patent without explicitly addressing its content. Maintain a structured and logical flow from basic recall to deeper conceptual understanding and practical application. Basic Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Text: {patent_summary} Generate a set of {qnum} questions that test a reader’s ability to recall fundamental concepts, key terms, and basic components necessary to understand the given patent. The questions should focus on defining terms, identifying parts, and recognizing the primary function of the described technology. Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }} Conceptual Level Question Prompt: You are given the following patent: Patent Title: {patent_title} Patent Text: {patent_summary} Generate a set of {qnum} questions that assess a reader’s comprehension of the given patent text by requiring them to explain how different elements of the invention work together. The questions should encourage the reader to summarize the design, describe interactions between components, and interpret the intended improvements of the invention. Output Format (JSON): Your response should be formatted as valid JSON: {{ “1”: “question1”, “2”: “question2”, “3”: “question3” }}
@@
Prompt for Answering Questions with Scientific Knowledge You are a scientific reasoning assistant. Use the following context to answer the question. Focus only on the information provided. Context: // retrieved scientific knowledge {context1} {context2} {context3} Question: {question} 1. Identify and summarize key points from the context relevant to the question. 2. Try your best to answer the question. Return your answer in the following JSON format: {{ “answer”: “your answer” //your answer to the question }}
@@
Prompt for the Splitting Task Judgment You are a patent judge. Below are two patent texts with extra information. Your task: judge whether they describe the same patent. Patent 1: {patent1} {qa_block1} Patent 2: {patent2} {qa_block2} Strictly return your concise answer in the following JSON format: {{ “label”: 1 // 1 if they describe the same patent, 0 if not “confidence”: 10 // your confidence score from 0 to 10, with 10 representing the highest confidence }} // Note: // This applies to the major experiments in the paper: "Do you think they are the same patent?" — when, in fact, the patents are judged to be *different* by human experts. // Invert this setup: Replace "same" with "different", and return 1 if the LLM concludes that the patents are *different*. This yields an alternative version of the experiment (See Section 3 Figure 2 top row in the paper).